Goldwing:

What is the exponential function modeled by the following table? a f(x) = 2x b f(x) = x2 c f(x) = x3 d f(x) = 3x

Goldwing:

Macho14:

is b x^2 and c x^3?? I am confused on how you wrote it

Macho14:

I will just assume unless you tell me otherwise

Goldwing:

Yes,it is;sorry about that.@macho14

Macho14:

oh ok thx I figured

Macho14:

how'd your last question go? did you get it right?

Goldwing:

Not quite sure yet but I think it is correct.How I don't know yet is due to waiting to get all my questions I have clarified , before submitting the assignment, you know...

Macho14:

yeah, this ones confusing, are you sure this is all the information, something seems off

Goldwing:

Did you see the screenshot provided ,yet.?

Macho14:

yeah, what do you think the answer is, wait, I see what I did wrong

Macho14:

So, first of all, we know the function f(x) is representative of y values

Goldwing:

Well,I think the answer is either between letter choice a or c?

Macho14:

hmm, none of these functions match any of the given points on the table

Macho14:

OOOOOOOH, thats the issue, it's 3^x, not 3x

Macho14:

duh, I should've known, since this is an exponential function

Macho14:

I just ruled out the first and last

Macho14:

thinking that they were wrong

Macho14:

b and c, are not the answer, so that leaves a and c

Goldwing:

Sorry about any confusions,@macho14 ,from now on I will try to add more clarity to the posts'.

Macho14:

yeah, it's no problem, I was just confused on how I turned stupid all of the sudden XD

Goldwing:

I can totally relate especially in Mathematics,it is one of the worst feelings,at least involving academics you know...Aside from that,did you break it down to the what the value would be as I as well agree that either a or c will be the value,etc.

Macho14:

I have the answer: the right answer is either a or c, plug in the x values on the table for answer a. \[2^x\] plug in first x value. \[2^3\]

Macho14:

it corresponds to the correct f(x) value

Macho14:

\[2^3=8\]\[2^4=16\]\[2^5=32\]

Macho14:

howeverAanswer a looks to be correct, but just to check, let's look at answer d: \[3^3\neq8\]

Macho14:

so \[f(x)=2^x\] must be correct

Macho14:

any questions, you can ask me, gl